Fermi National Laboratory
 

Accelerator Update for May 15, 2009 - May 18, 2009

 

The Integrated Luminosity 5/11/09 to 5/18/09 was 58.04 pb-1.

Friday May 15

The day shift began with Operations monitoring the Tevatron (TeV) store 7054 (12.39E31) and the Recycler stash (77.7E10), with the Antiproton Source stacking (26.4E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.

At 2:30 PM, the Pbar Debuncher injection kicker stopped pulsing. Technicians repaired a cable on the high voltage power supply.

Pbar resumed stacking at 4:15 PM.

Operations terminated store 7054 at 11 PM and then put the TeV into HEP shot setup.

Saturday May 16

The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (260.1E10) and the stack (42.8E10), and with NuMI and MiniBooNE taking beam.

Operations established store 7055 by 12:45 AM with an average initial luminosity of 31.8E31.

Operations established beam to MTest at 2:15 AM.

Operations terminated store 7055 at 3 PM and put the TeV into HEP shot setup.

Operations established store 7056 at 4:30 PM with an average initial luminosity of 31.8E31.

Control issues with a M01 LCW system occasionally inhibited spill to MTest. Instrumentation personnel were called in to investigate.

Operations disabled the MTest beam at 8 PM.

Sunday May 17

The midnight shift began with Operations monitoring store 7056 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

Operations established beam to MTest at 2:10 AM.

Operations terminated store 7056 at 6:55 AM and put the TeV into HEP shot setup at 7:05 AM.

Operations established store 7057 at 8:35 AM with an average initial luminosity of 32.7E31.

A MI Lambertson magnet (LAM52) tripped off on a magnet temperature indication. Operations contacted EE Support personnel to investigate. They accessed MI and reterminated a lead wire for one of the magnet’s Klixons. During the MI turn-on, operators found a ground fault between MI-40 and MI-60. Operations called in MI personnel. The found the ground fault on a quadrupole magnet at the 428 location. Operations called in Mechanical Support personnel to replace the faulty magnet.

Monday May 18

The midnight shift began with Operations monitoring store 7057 and the stack, with MI in access to replace a Lambertson magnet, and with NuMI and MiniBooNE off for the MI access.

The MI QF428 magnet replacement lasted throughout the midnight shift.

The Plans for Monday

The plans for today are to finish the MI Lambertson installation, which should occur around noon today. D0 experimenters reported a problem that had stopped them from taking data. Repairs would require them to access and open up their detector, which would take around nine hours. The Run Coordinator granted them their access starting at 9 AM this morning.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 05/18/2009    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory